OpenStudy (anonymous):

the formula A= P(1.09)^t is an example of exponential growth with base 1.09. Determine an equivalent continuous growth formula using the base e, A=Pe^kt. Can someone explain the answer in detail? I know the final answer, but I don't understand the process of getting there

OpenStudy (anonymous):

dy/dt = 0.09 y 9% growth per time period dt dy/y = 0.09 dt ln(y) = 0.09 t A = P exp(0.09 t) continuous compounding rather than yearly compare at t=1 year A/P = 1.09 vs A/P = exp(0.09) = 1.094

OpenStudy (anonymous):

I don't understand? the answer is A= Pe^0.086t?